### v2

# Binding energies, charge radii, spins and moments: odd-odd Ag isotopes and discovery of a new isomer

Abstract
We report on the masses and hyperfine structure of ground and isomeric states in 114,116,118,120Ag isotopes, measured with the phase-imaging ion-cyclotron-resonance technique (PI-ICR) with the JYFLTRAP mass spectrometer and the collinear laser spectroscopy beamline at the Ion Guide Isotope Separator On-Line (IGISOL) facility, Jyväskylä, Finland. We measured the masses and excitation energies, electromagnetic moments, and charge radii, and firmly established the nuclear spins of the long-lived states. A new isomer was discovered in 118Ag and the half-lives of 118Ag long-lived states were reevaluated. We unambiguously pinned down the level ordering of all long-lived states, placing the inversion of the and states at (). Lastly, we compared the electromagnetic moments of each state to empirical single-particle moments to identify the dominant configuration where possible.
I Introduction
Hyperfine structure data and mass measurements on the ground and long-lived isomeric states provide information on their single-particle and collective nuclear properties Yang et al. (2023); Yamaguchi et al. (2021). The former is sensitive to the electromagnetic properties of the nucleus, i.e. magnetic dipole and electric quadrupole moments, charge radii, and nuclear spins Campbell et al. (2016). The latter can provide excitation energies and masses, and thus the level ordering of the isotope Dilling et al. (2018). Together, they give insight into the nuclear configuration while also probing its collective behavior and therefore the degree of configuration mixing. These observables allow identification and investigation of structural changes, e.g. (sub)shell closures or shape changes throughout an isotopic chain Otsuka et al. (2020).
The region below tin () is of high interest as it allows for probing shell evolution below the closed proton shell at and between the neutron shell closures at . While radioactive isotopes between palladium () and tin () have been studied before via mass measurements Hager et al. (2007); Breitenfeldt et al. (2010); Hakala et al. (2012); Kankainen et al. (2013); Atanasov et al. (2015); Babcock et al. (2018); Nesterenko et al. (2020); Manea et al. (2020); Izzo et al. (2021); Mougeot et al. (2021); Nesterenko et al. (2023); Nies et al. (2023); Jaries et al. (2023); Ge et al. (2024); Jaries et al. (2024); Ruotsalainen et al. (2024) and laser spectroscopy Yordanov et al. (2013); Ferrer et al. (2014); Hammen et al. (2018); Gorges et al. (2019); Yordanov et al. (2020); Reponen et al. (2021); Geldhof et al. (2022); Vernon et al. (2022); Karthein et al. (2024); de Groote et al. (2024), data on neutron-rich odd-odd Ag isotopes is absent or incomplete.
The masses of neutron-rich silver isotopes have previously been studied using the time-of-flight ion-cyclotron-resonance (ToF-ICR) technique at ISOLTRAP, CERN Breitenfeldt et al. (2010) but the long-lived states could not be separated due to the limited resolving power of the ToF-ICR method. The phase-imaging ion-cyclotron-resonance (PI-ICR) Eliseev et al. (2013, 2014); Nesterenko et al. (2018) technique has a much higher resolving power enabling the measurements of the long-lived states as has already been demonstrated with the even-A silver isotopes in Ref. de Groote et al. (2024).
Multiple decay spectroscopy studies have been performed on neutron-rich odd-odd Ag isotopes identifying energies and multipolarities of the internal isomeric transitions Koponen et al. (1989); Batchelder et al. (2005, 2012). However, they were conducted using beams without isomeric or even isobaric purification. As a result, spin assignments, which are based on logft-values and branching ratios Janas et al. (1993); Wang et al. (2003); Batchelder et al. (2005, 2012), are often tentative, sometimes causing conflicting conclusions between different publications Koponen et al. (1989); Janas et al. (1993); Wang et al. (2003); Rissanen et al. (2007).
The combination of mass and laser spectroscopy measurements at the same facility with the same production method can remedy this situation by providing a complementary insight into the studied isotope, as these techniques are sensitive to different observables. Furthermore, states and excitation energies can be assigned to their spin-parities unambiguously by comparing the yields between the two experiments as demonstrated earlier in Ref. de Groote et al. (2024).
Recent mass spectrometry and collinear laser spectroscopy results de Groote et al. (2024) have been reported on neutron-rich odd-even 113-121Ag, unambiguously pinning down the spins and excitation energies and showing the continuation of the and spins for these isotopes. Notably, the g-factors of the states show little variation for varying neutron numbers, indicating a relatively pure three-hole (g configuration, while for the states they show a clear decreasing trend indicating a varying degree of configuration mixing between different isotopes.
This work presents mass and collinear laser spectroscopy measurements performed in parallel on 114,116,118,120Ag. The results are combined with the current decay data to firmly establish spin and parity assignments as well as the level ordering for three long-lived states observed in 116,118,120Ag. The electromagnetic moments are compared to empirical single-particle moments and a dominant configuration is attributed where possible. We already reported on the mean-squared charge radii of the lowest-spin states of 114-120Ag in a previous publication Reponen et al. (2021). Here, we re-investigated these radii resulting in larger uncertainties and we provide the electromagnetic moments, masses and excitation energies, and the nuclear mean-squared charge radii for all long-lived states in this mass range.
II Experimental methods
Radioactive silver isotopes were produced at the IGISOL facility by impinging a 25 MeV proton beam onto a thin uranium foil, where the protons induced the uranium nuclei to fission. Reaction products were stopped in a helium-filled gas cell, operated at 300 mbar of gas pressure. The gas flow guided these ions into a sextupole ion guide Karvonen et al. (2008), after which they were accelerated to 30 kV. The ions were then mass-separated using a 55-degree dipole magnet and injected into a helium-filled radiofrequency quadrupole (RFQ) Nieminen et al. (2001). Here, the ions were buffer-gas cooled, bunched, and transported to either the JYFLTRAP double Penning trap mass spectrometer Eronen et al. (2012) or the collinear laser spectroscopy beamline de Groote et al. (2020). The storage time of the ions in the RFQ varied depending on the measurement cycle in the mass measurements and was fixed at 100 ms for the laser spectroscopy measurements.
At JYFLTRAP, the ions were first cooled and centered on the trap axis by using a mass-selective buffer-gas cooling technique Savard et al. (1991) in the preparation trap, and then transferred to the measurement trap through a 1.5-mm diaphragm. After a few ms, the ions of interest were transferred back to the preparation trap for an additional step of cooling. Two steps were needed for effective cooling and centering of the ions of interest due to the large amount of contaminant ions. For 114Ag+ ions, a Ramsey-cleaning method Eronen et al. (2008) with an excitation pattern 5-30-5 ms was additionally applied to remove neighboring isobars. Finally, the ions of interest with charge-to-mass ratio were sent to the measurement trap, where their cyclotron frequency in the magnetic field was determined using the PI-ICR technique Eliseev et al. (2013, 2014); Nesterenko et al. (2018).
The phase accumulation time () in the PI-ICR method was chosen in such a way as to separate the isomeric states in the silver isotopes and to ensure that the projection of the cyclotron motion onto the detector does not overlap with any possible isobaric or molecular contamination. The was 400 ms for 114,118,120Ag and their respective isomeric states while for 116Ag it was 600 ms. The measurements were performed for singly-charged ions. The magnetic field strength was determined using either 133Cs (mass excess keV Wang et al. (2021)) produced in the offline surface ion source Vilén et al. (2020) or the isobaric species produced in fission as the reference. The atomic mass was determined from the cyclotron frequency ratio of the reference ion to the ion of interest as
where and are the atomic mass of the reference ion and the electron mass, respectively. The binding energy of the missing electron was neglected. The systematic uncertainties were included in the final uncertainty of the cyclotron frequency ratios Nesterenko et al. (2021). The count-rate class analysis Roux et al. (2013) was performed for the frequency ratios to take into account ion-ion interactions in the measurement trap.
In the laser spectroscopy line, the ions were first neutralized using a charge exchange cell filled with potassium. This cell was heated to about 130 degrees Celsius using a NiCr heating wire wrapped around the center, while the ends were kept at 70 degrees Celsius using a liquid cooling loop filled with GALDEN heat transfer liquid. By changing an acceleration potential applied to the cell, the atoms were Doppler-shifted into resonance with the counter-propagating continuous-wave 328 nm laser beam to probe the transition. The laser wavelength was changed in order to keep the scanning voltage between 0 and 2 kV. Measurements were spread over three different experimental periods. For the first experimental run, the laser light was produced using an intra-cavity frequency-doubled Spectra Physics 380 dye laser which was pumped by a 5 W Verdi 532 nm laser. The other measurements used a Matisse Dye laser pumped by a 10 W Millennia 532 nm laser. Instead of intra-cavity doubling, an external WaveTrain frequency-doubler was used. By recording the number of photon counts observed by the photo-multipliers as a function of the wavelength in the rest frame of the atoms, the hyperfine structures of the Ag isotopes could be recorded. To reduce the background from scattered laser light, the recorded number of photons was time-gated such that only the photons measured when the atom bunch was in front of the photo-multiplier were used in the hyperfine spectra. Regular reference measurements were performed on 109Ag by injecting stable beams produced by a dedicated offline spark-discharge ion source Vilén et al. (2020) into the cooler-buncher. Repeated measurements of the same isotopes were performed across all three experiments as an additional systematic check; all three datasets were found to be in mutual agreement. The hyperfine (HF) spectra were fitted with the SATLAS analysis suite Gins et al. (2018).
The magnetic dipole moment and spectroscopic electric quadrupole moment were determined with Eqs. 2 and 3.
Here, is the nuclear spin and and are the hyperfine - and -constants. The and denote a non-optical reference measurement of the dipole and quadrupole moment from literature, and the and denote high-precision measurements of the hyperfine - and -constants from literature.
The changes in mean-squared charge radii are calculated using Eq. 4.
Here, is the isotope shift between isotopes A and which is extracted from the hyperfine spectra, is the change in mean-squared charge radius between these isotopes, m and are their masses, F and M are the field- and mass-shift atomic factors coming from atomic or empirical calculations and K (=0.976) is a correction factor to the field-shift factor for higher-order radial moments as used in Ref. Reponen et al. (2021); Geldhof et al. (2022). As seen in Eq. 4, the factor is the scaling constant for the contribution due to the difference in nuclear volume to the isotope shift, while the factor is the scaling constant for the contribution due to the difference in nuclear mass between isotopes A and .
In the case that the HF structure of multiple nuclear states was overlapped, the highest peak in the HF spectrum was associated with the highest spin nuclear state, as this is the most produced in this experiment Cannarozzo et al. (2023). Moreover, as the number of peaks in the HF structure is larger than the number of free variables (A, A, B and isotope shift ) in 116,118,120Ag, there is no ambiguity in assigning resonances to the difference nuclear states.
In this work, the statistical uncertainties originate from the uncertainty from the fit and the scan-to-scan variation of each fit parameter. The isotope shifts from the lowest spin states, reported in Ref. Reponen et al. (2021), are reported here again, including the scan-to-scan variation. This resulted in consistent values, but larger uncertainties. The systematic uncertainties on the nuclear observables originate from the reference moments and hyperfine - or -constants for the dipole and quadrupole moments, and the field- and mass-shift factor and nuclear masses for the changes in mean-squared charge radii. While the hyperfine anomaly in silver can go up to a few percent Persson (2013), we decided to neglect this contribution as it is not known for the studied isotopes and requires further development. The dipole moment is calculated with a reference = MHz Dahmen and Penselin (1967), = Sahm and Schwenk (1974); Mertzimekis et al. (2016) assuming a zero hyperfine anomaly, and the quadrupole moment with = MHz Fischer et al. (1975) and = b Berkes et al. (1984). The changes in mean-squared charge radii are extracted with newly-calculated atomic field- and mass-shift factors, F = MHz/fm2 and M = GHz u Ohayon et al. (2024), different from the ones reported in Ref. Reponen et al. (2021) based on empirical estimates.
III Results and Discussion
All the results are summarized in Tables 1 and 2. The masses, spins, and state ordering are discussed in Sec. III.1, the dipole moments and configurations in Sec. III.2, and the quadrupole moments and charge radii in Sec. III.3.
III.1 Level ordering and spin assignments
The following section discusses the studied silver isotopes’ spin assignments and level ordering. Often, spin assignments can be made using laser spectroscopy data based on the ratio of - or -constants of the atomic ground and excited states. However, in this case, the ratio is rather small (1/53.5), which when combined with the size of the experimental uncertainties, prevents such assignments. Furthermore, the atomic ground state is known to exhibit a large hyperfine anomaly Persson (2013), further complicating these assignments based on this ratio. We opted to determine the spins in this work for each state separately based on trends in the nuclear observables (e.g. changes in mean-squared charge radii), combined with existing decay spectroscopy data.
III.1.1 114Ag
One long-lived nuclear state, the ground state is known in 114Ag. This state is observed in the mass and laser spectroscopy measurement, the latter indicated in Fig. 1. We note that the known isomer Penttilä et al. (1990) has a half-life of 1.50(5) ms Kondev et al. (2021) which is much shorter than our measurement cycles thus it was not present in the beam. The mass excess measured in this work differs from the AME2020 value Wang et al. (2021), which is based on the ISOLTRAP measurement Breitenfeldt et al. (2010), by 1.6 and it is 4 times more precise.
The observed hyperfine structure is consistent with the literature assignment of Koponen et al. (1989) as only two resonances could be observed on the high-frequency side. A nuclear spin would involve three resonances with approximately equal intensities. Furthermore, analysis assuming resulted in a large isotope shift ( MHz), which would in turn result in an unphysical change in mean-squared charge radius. Our data therefore firmly confirms the spin of the ground state of 114Ag to be .
| Nuclide | Ref. | (keV) | (keV) | () (keV) | (keV) | (keV) | |||
| 114Ag | 4.6(1) s | 133Cs | 0.857 066 064(10) | ||||||
| 116Ag | 3.83(8) m | 116Agm2 | 0.999 998 740(38) | ||||||
| 116Agm1 | 20(1) s | 116Agm2 | 0.999 999 198(27) | 49.5(51) | 47.9(1) | ||||
| 116Agm2 | 9.3(3) s | 133Cs | 0.872 134 683(25) | 136.0(41) | 129.80(22) | ||||
| 118Ag | 1.92(9)1 s | 118Agx | 0.999 998 876(11) | 0 | 127.63(10) | ||||
| 118Agx | 5-61 s | 133Cs | 0.887 207 238(22) | 123.5(12) | 0 | ||||
| 1.92(9)1 s | – | ||||||||
| 120Ag | 1.52(7) s | 133Cs | 0.902 285 837(22) | ||||||
| 120Agm1 | 0.94(10) s | 120Ag | 1.000 001 030(31) | 115.1(35) | 0(50)# | ||||
| 120Agm2 | 384(22) ms | 120Ag | 1.000 001 817(36) | 203.0(40) | 203.0(2) |
This work
| Nuclide | A(S1/2) | A(P3/2) | B(P3/2) | Q | r | |||
| (MHz) | () | (MHz) | (MHz) | (b) | (MHz) | (fm2) | ||
| 114Ag | 19215(22) | 352(4) | 59(3) | 0.404(5)[6] | ||||
| 116Ag | 805(43) | 15.1(8)1 | 0.522(3)[8] | |||||
| 116Agm1 | 5329(5) | 96.6(10) | 234(38) | 0.494(8)[7] | ||||
| 116Agm2 | 3304.0(5) | 60.10(12) | 293(7) | 0.490(9)[7] | ||||
| 118Ag | 5911(2) | 106.0(12) | 297(13) | 0.618(6)[9] | ||||
| 118Agm1 | 0 | 0 | 0 | 0 | 0 | 0.632(4)[9] | ||
| 118Agm2 | 3812.9(10) | 68.8(4) | 408(6) | 0.612(4)[9] | ||||
| 120Ag | 6045(4) | 107.9(16) | 299(17) | 0.7067(12)[103] | ||||
| 120Agm1 | 0 | 0 | 0 | 0 | 0 | 0.7317(17)[106] | ||
| 120Agm2 | 3651(3) | 65.5(9) | 371(7) | 0.712(2)[10] |
The A/A was fixed to the ratio of 109Ag as the hyperfine structure was unresolved in our data.
III.1.2 116Ag
Three long-lived states in 116Ag are known in literature Kondev et al. (2021). Spins and were assigned from -decay spectroscopy Batchelder et al. (2005). All three states were observed and resolved in the mass and laser spectroscopy measurements.
The mass-excess value of the 116Ag ground state determined at JYFLTRAP differs by 2.4 from the AME2020 value Wang et al. (2021), which is based on the ISOLTRAP measurement Breitenfeldt et al. (2010). The measured excitation energy of 116Agm1 is in a good agreement with the NUBASE2020 value while the 116Agm2 excitation energy differs by 2.8 Kondev et al. (2021). It should be noted that by using the mass excess of 116Agm2 from this work and the excitation energy from the NUBASE2020 evaluation Kondev et al. (2021), the calculated mass-excess value of the 116Ag ground state ( keV) is in an agreement with AME2020 value.
Fig. 1 shows the hyperfine structures measured for 116Ag in this work. Two states exhibit rather large hyperfine splitting, while the third is smaller. Crucially, none of them show only one peak, as would be expected for an state. We can thus rule out this spin assignment for the ground state. Therefore, knowing that the long-lived states are connected via E3 transitions Batchelder et al. (2005), we can infer that the and assignments are also incorrect.
While our data does not provide a clear preference for either spin or for the ground state, the decay spectroscopy strongly favors the assignment. In the case of , a substantial feeding via allowed -decays to the states in the daughter 116Cd isotope can be expected, however, it has not been observed in the decay spectroscopy study of 116Ag Batchelder et al. (2009). In addition, a would imply that the second isomeric state is due to the presence of the E3 transitions, but in Refs. Batchelder et al. (2005, 2009) feeding to neither 8+ nor (9-) states in 116Cd was observed. This firmly fixes the spins to , and for the ground and two isomeric states respectively. We note that these assignments resolve the problem of two missing transitions between proposed (8-) and (5+) states, as reported in Refs. Porquet et al. (2003); Kim et al. (2017).
III.1.3 118Ag
Two long-lived states in 118Ag are known in literature Kondev et al. (2021) and this was also observed by the mass measurement, as shown in Fig. 2. However, three states were unambiguously identified in the collinear laser spectroscopy experiment. The phase-accumulation time in the PI-ICR measurement was varied from 50 ms to 800 ms, and the final measurements were performed with ms. Most likely the third state, not observed in the PI-ICR measurement, lies close to one of the observed states with a mass difference keV (based on the accumulation time). The excitation energy of 118Agx determined at JYFLTRAP deviates by 4.1(12) keV from the precisely-known NUBASE2020 value of 127.63(10) keV Kondev et al. (2021), based on the E3 -decay observed in Refs. Fogelberg et al. (1971); Koponen et al. (1989); Rissanen et al. (2007). As this difference is negative, the third state must be lower in energy than 127.63(10) keV, thus we conclude that the 127.63(10) keV E3 transition connects the ground state and the second isomeric state.
To determine which two states are unresolved in the PI-ICR spectra, we compare the production of each isomer between the mass and laser spectroscopy measurements. In the former, the ion ratio between the ground state and the isomeric spot is about 0.46. The decay losses in the trap were neglected as the phase accumulation time is much shorter than the half-lives of the long-lived states in 118Ag. In the latter, we calculated the integral of hyperfine peaks of a specific isomer to estimate its relative production rate. Eight different combinations of peaks were considered to verify the consistency of these rates, see Supplementary Material VI. The extracted rate of the low-, medium-, and high-spin states is 10.2(4)%, 30.7(11)%, and 59.1(13)%, respectively. These two results are compatible with each other only when the medium-spin state is the ground state and the low- and high-spin states are the excited states, since .
From the laser spectroscopy data, we can firmly conclude that the low-spin state is a state, as only a single peak is observed in the hyperfine structure, shown in Fig. 1. The parity is inferred from the only possible configuration which can yield a spin zero in this region of the chart, coupling a proton in the p1/2 and a neutron in the s1/2 orbital. Additionally, we conclude that the medium-spin state is an state, as a spin 3 or 5 would yield of 0.7069(55) or 0.5291(51) fm2 respectively, which are very far from the trend shown in Fig. 3. Consequently, the E3 transition observed in literature must be between this state and the high-spin state, as an E3 -transition cannot connect the and the states. This fixes the high-spin state to , where the parity is inferred as this can only be reproduced by coupling a proton in the g9/2 and a neutron in the h11/2 orbital. Thus, the state must be positive parity.
Based on the combined collinear laser spectroscopy and mass measurements, we firmly assign the ground state as , the first isomer as , and the second isomer as . Using the precisely determined -ray energy from Kondev et al. (2021), the state is at 127.63(10) keV, while the excitation energy of the isomer can be estimated to be between 103-123 keV. The lower limit is determined by the excitation energy of the second isomer and the resolution of the PI-ICR measurement, i.e. keV, while the upper limit is the excitation energy of 118Agx from the PI-ICR measurement, 123.5(12) keV. Similarly to 116Ag, the observation of the isomeric state allows for the removal of two suggested but unobserved transitions from the excited states scheme in 118Ag Wang et al. (2017); Kim et al. (2017).
In addition to the hyperfine structure scans of the and states, multiple collinear laser spectroscopy measurements on these states have been performed with cooler-buncher cooling times, ranging from 100 ms up to 1 s. This should, in principle, allow for half-life extraction. However, changes in experimental conditions, such as primary beam current, and transmission losses also influence the number of atoms, so we opted for a relative measurement to cancel out these changes. The ratio of relative yields , normalized for acquisition time, is proportional to the activity ratio . By analyzing the change of these ratios as a function of cooling time, as is done in Fig. 4, one can extract the difference between decay constants of the () and the () state:
where is a normalization constant and . The fitted difference is 4(4) /s, i.e. consistent with the two states having the same half-life within our experimental uncertainties.
While the existence of the state was unknown prior to this work, its half-life can be extracted based on the previous measurements of the 127.63(10) keV -ray. Two values have been reported in the literature, 1.9(1) s in Ref. Koponen et al. (1989) and 2.0(2) s in Ref. Hill (1978). We take their weighted average, 1.92(9) s as the half-life of the state. Based on these results and our measurement, we extracted the half-life of the state to be 1.92(9) s. The state will likely have a half-life of the order of 5-6 s as reported in Refs. Fritzs and Griffiths (1967); Weiss et al. (1968), rather than the more recently reported 3.7(2) s and 3.76(15) s Hill (1978); Koponen et al. (1989) which were measured for isomeric mixtures.
The discovery of a new isomer and the new order of the long-lived states imply a different interpretation of the two decay spectroscopy studies of 118Ag Wang et al. (2003) and 118Pd Janas et al. (1993) where only two long-lived states in 118Ag with different spins and a different level order were proposed. In both of these publications, the production method was the same as in this work and the beam contained the entire isobar, thus, all three long-lived states of 118Ag were present. As a result, these decay schemes should be reevaluated. In particular, as the half-lives of the and the states determined in this work are identical, the decays of these two states could not be separated in Ref. Wang et al. (2003). Additionally, in Ref. Janas et al. (1993), there was no sensitivity to distinguish between the decay of 118Pd and the internal transition decay of 118Agm2 as the half-lives of 118Pd (T1/2 = 1.9(1) s) and 118Ag() (T1/2 = 1.92(9) s) are identical within experimental uncertainties. Consequently, the population of the 127.63(10) keV state in the -decay of 118Pd remains doubtful.
III.1.4 120Ag
Three long-lived states in 120Ag are known in literature Kondev et al. (2021), all of which are observed in the mass and laser spectroscopy measurements, see Figs. 1 and 5. Spin-parities and were previously assigned based on the -feeding to the states in Cd and systematics in Ref. Batchelder et al. (2012). However, the relative order of the low- and the medium-spin states remained unknown.
The energy difference between 120Aggs and 120Agm2 measured in this work, 203.0(40) keV, matches the known energy difference between the and state of 203.0(2) keV Kondev et al. (2021). Therefore, we conclude that the observed isomer at 115.1(35) keV is the low-spin state and its excitation energy was determined for the first time. The mass-excess value of 120Ag determined at JYFLTRAP differs from the AME2020 value Wang et al. (2021) by 5 . However, it is exclusively based on the ISOLTRAP value Breitenfeldt et al. (2010) where the three long-lived states were not resolved.
Our hyperfine spectra feature only a single peak for the low-spin isomer of 120Ag, as is the case with 118Ag. This means that there is no hyperfine splitting present which is only possible for a state or a state with a near-zero dipole and quadrupole moment. However, for a state, conversion electrons at 110-115 keV would be expected in Ref. Batchelder et al. (2012), but none are seen in their spectra. Thus, we firmly assign a to this state, where the parity is inferred as this spin can only be reproduced by coupling a proton in the p1/2 and a neutron in the s1/2 orbital.
In Ref. Batchelder et al. (2012), the absence of -feeding to the state and the observation of -feeding to in the daughter 120Cd led to the assignment for the high-spin state. This state decays with an E3 -transition to the medium-spin state which is then an state. The changes in the mean-squared charge radii determined in this work support this assignment. Analysis assuming an or yield of 0.7939(12) or 0.6112(10) fm2 respectively, which are very far from the trend shown in Fig. 3. Therefore we conclude that the and are the spins for these two states. The parities are assigned with certainty as the can only be reproduced by coupling a proton in the g9/2 and a neutron in the h11/2 orbital. We note that the ordering in 120Ag is the same as in 118Ag thus our measurement firmly places the state order change at .
III.2 Magnetic dipole moments and configuration
The following section discusses the magnetic dipole moments and configuration of the different silver isotopes measured in this work. Empirical single-particle moments are calculated with the addition rule from Ref. Heyde (1994), given by:
In this formula, we take the dipole moment of a neighboring odd-, even- isotope for , and of a neighboring even-, odd- isotope for to calculate the empirical single-particle dipole moment of the odd-, odd- configuration.
III.2.1 106-114Ag
The -factors of the states follow a relatively constant trend across the isotopic chain, slightly decreasing at 114Ag as seen in Fig. 6. The g-factors of the states are in reasonable agreement with the calculated moments using and - which can be interpreted as the and the configuration. While the magnitude and trend of the empirical single-particle moments are similar for both configurations, previous literature suggests a dominating configuration Fischer et al. (1975); Winnacker et al. (1976); Golovko et al. (2010). In Ref. Ferrer et al. (2014) an evolution from a single-proton hole in the orbital to a three-proton hole in the orbital from is reported for the high-spin states of these isotopes.
III.2.2 116-120Ag
The small g-factor of 0.107(6) for the state of 116Ag shows an indication that the g9/2 and h11/2 orbitals are unlikely to play a role in this state as its empirical single-particle moment InCd and single-particle moment () differs strongly from the experimental g-factor. The best-matching empirical single-particle moment is the AgCd, which is also not in good agreement with the experimental g-factor. The configurations for which the single-particle moments agree best are p(d and p(g. We conclude that the configuration of this state is very mixed.
Fig. 7 compares the states with the results of the additivity rule using several empirical moments. A proton in the (g (Ag(7/2)) configuration coupled with a neutron in the single-particle g7/2 (Sn(7/2), a neutron in the mixed Sn, or a neutron in the mixed Cd configuration result in similar agreement with the experimental g-factors as seen in Fig. 7). Likely, there is a high amount of mixing in the neutron configurations, which is expected due to the high density of positive parity orbitals between N = 50 and N = 82.
Lastly, the isomers shown in Fig. 8 are in relatively good agreement with the empirical single-particle moments, as expected since fewer orbitals can couple to a state and thus less configuration mixing is possible. The closest-matching empirical moments are the and the , interpreted as the gh11/2 and (gPd configurations, both of which yield similar agreement with the data. However, no dominating configuration for either proton or neutron can be deduced from the g-factors.
III.3 Charge radii and electric quadrupole moments
The following section discusses the charge radii and spectroscopic quadrupole moments of the different silver isotopes measured in this work. The charge radii are compared to neighboring chains and the quadrupole moments to empirical single-particle moments. The empirical single-particle moments are calculated with the addition rule from Ref. Heyde (1994), given by:
In this formula, we take the quadrupole moment of a neighboring odd-, even- isotope for , and of a neighboring even-, odd- isotope for to calculate the empirical single-particle quadrupole moment of the odd-, odd- configuration, similar to Eq. 6. The and indicate the Wigner 3-j and Wigner 6-j symbols respectively.
Fig. 9 shows the absolute charge radii of Pd () up to Sn () with two different sets of F and M factors for the Ag isotopic chain. The empirical F and M factors Reponen et al. (2021) are MHz/fm2 and GHz u, and the analytical-response relativistic coupled-cluster (AR-RCC) F and M factors from Ohayon et al. (2024) are MHz/fm2 and GHz u. There is a 3 difference in the field shift and a 2 difference in mass shift factors between the empirical and ab-initio calculation. In Fig. 9, only a small difference can be seen between the two trends on the neutron-rich side. Note that in Ref. Reponen et al. (2021) the charge radii of the lowest-spin states were reported, while here in Fig. 9 the charge radii for the ground states are shown.
The spectroscopic quadrupole moments for odd-odd silver are shown in Fig. 10. The and states show a positive spectroscopic quadrupole moment similar to the indium Vernon et al. (2022); Karthein et al. (2024) and odd-A silver de Groote et al. (2024) isotopes. The states do not agree very well with the empirical quadrupole moments calculated with the unpaired neutron in Cd or Sn, which show a linear increase with increasing as in odd-A Cd isotopes Yordanov et al. (2013). However, there is an excellent agreement with the empirical moment for , interpreted as the Ag(gPd configuration. This is in contrast with the trend for the g-factors in Fig. 8 which shows a relatively good agreement for both configurations.
The empirically calculated quadrupole moments for the states agree relatively well with the experimentally measured moments. However, there seems to be no preference for a single- or three-hole proton configuration in the g9/2 orbital, similar to the state. A more detailed interpretation of these moments would require e.g. large-scale shell model calculations.
IV Conclusion
The odd-odd isotopes 114-120Ag have been studied with collinear laser spectroscopy and PI-ICR mass spectrometry at IGISOL in Jyväskylä, Finland. All spins, parities, electromagnetic moments, and level ordering of the long-lived states for 114-120Ag are determined unambiguously. Specifically, the spins of 116Ag have been reassigned. Furthermore, a new isomer has been identified in 118Ag, whose presence reinterprets the level ordering, half-lives, and spins previously suggested in the literature. The existing decay scheme of 118Ag has been shown to be incorrect. We thus identify a need to perform decay spectroscopy of isomerically purified beams in a future experiment. Lastly, the previously unknown excitation energy of the low-spin isomer in 120Ag has been measured.
The charge radii were found to vary smoothly with N, similar to the neighboring elements. From the magnetic dipole and quadrupole moments, it can be concluded that these isotopes have a rather mixed neutron configuration, in line with the more collective properties of Pd. This illustrates that silver sits at the transition between the spherical Sn isotopes, and the well-deformed isotopic chains .
Multiple campaigns are ongoing to understand this region further. A decay spectroscopy experiment has been performed at IGISOL, Jyväskylä to reassign the -transitions in the decay of 118Pd. An experimental campaign at CRIS at ISOLDE/CERN is ongoing to perform laser spectroscopy on neutron-rich silver, approaching the magic shell closure de Groote et al. (2020). Moreover, a muonic X-ray experiment on 107,108m,109Ag at PSI is being prepared to establish high-precision field and mass shift factors Cocolios (2020).
More work is also being done from the theoretical perspective. Atomic ab initio calculations are ongoing to calculate the hyperfine and constants in the 5s , 5p and 5p atomic levels and the field and mass shift factors in the 4d105s 4d105p and 4d105s 4d105p transitions. This is calculated in parallel with new state-of-the-art DFT calculations for the diamagnetic and paramagnetic shielding constants to revise the reference magnetic dipole moment from Ref. Sahm and Schwenk (1974).
This work provides a step towards understanding the nuclear structure in the largely unexplored region moving away from the spherical Sn isotopes.
Acknowledgements.
RPDG received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 844829. The funding from the European Union’s Horizon 2020 research and innovation program under grant agreement No. 771036 (ERC CoG MAIDEN) and Research Council of Finland (Grant Nos. 314733, 320062, 318043, 295207, 327629 and 354968) are gratefully acknowledged. This work was supported by the UK Science and Technology Facilities Council (STFC). We acknowledge W. Nörtershäuser for the use of the charge-exchange cell.References
- Yang et al. (2023) X. Yang, S. Wang, S. Wilkins, and R. G. Ruiz, Progress in Particle and Nuclear Physics 129, 104005 (2023).
- Yamaguchi et al. (2021) T. Yamaguchi, H. Koura, Y. A. Litvinov, and M. Wang, Progress in Particle and Nuclear Physics 120, 103882 (2021).
- Campbell et al. (2016) P. Campbell, I. Moore, and M. Pearson, Progress in Particle and Nuclear Physics 86, 127 (2016).
- Dilling et al. (2018) J. Dilling, K. Blaum, M. Brodeur, and S. Eliseev, Annual Review of Nuclear and Particle Science 68, 45 (2018).
- Otsuka et al. (2020) T. Otsuka, A. Gade, O. Sorlin, T. Suzuki, and Y. Utsuno, Rev. Mod. Phys. 92, 015002 (2020).
- Hager et al. (2007) U. Hager, V.-V. Elomaa, T. Eronen, J. Hakala, A. Jokinen, A. Kankainen, S. Rahaman, S. Rinta-Antila, A. Saastamoinen, T. Sonoda, and J. Äystö, Phys. Rev. C 75, 064302 (2007).
- Breitenfeldt et al. (2010) M. Breitenfeldt, C. Borgmann, G. Audi, S. Baruah, D. Beck, K. Blaum, C. Böhm, R. B. Cakirli, R. F. Casten, P. Delahaye, M. Dworschak, S. George, F. Herfurth, A. Herlert, A. Kellerbauer, M. Kowalska, D. Lunney, E. Minaya-Ramirez, S. Naimi, D. Neidherr, M. Rosenbusch, R. Savreux, S. Schwarz, L. Schweikhard, and C. Yazidjian, Phys. Rev. C 81, 034313 (2010).
- Hakala et al. (2012) J. Hakala, J. Dobaczewski, D. Gorelov, T. Eronen, A. Jokinen, A. Kankainen, V. S. Kolhinen, M. Kortelainen, I. D. Moore, H. Penttilä, S. Rinta-Antila, J. Rissanen, A. Saastamoinen, V. Sonnenschein, and J. Äystö, Phys. Rev. Lett. 109, 032501 (2012).
- Kankainen et al. (2013) A. Kankainen, J. Hakala, T. Eronen, D. Gorelov, A. Jokinen, V. S. Kolhinen, I. D. Moore, H. Penttilä, S. Rinta-Antila, J. Rissanen, A. Saastamoinen, V. Sonnenschein, and J. Äystö, Phys. Rev. C 87, 024307 (2013).
- Atanasov et al. (2015) D. Atanasov, P. Ascher, K. Blaum, R. B. Cakirli, T. E. Cocolios, S. George, S. Goriely, F. Herfurth, H.-T. Janka, O. Just, M. Kowalska, S. Kreim, D. Kisler, Y. A. Litvinov, D. Lunney, V. Manea, D. Neidherr, M. Rosenbusch, L. Schweikhard, A. Welker, F. Wienholtz, R. N. Wolf, and K. Zuber, Phys. Rev. Lett. 115, 232501 (2015).
- Babcock et al. (2018) C. Babcock, R. Klawitter, E. Leistenschneider, D. Lascar, B. R. Barquest, A. Finlay, M. Foster, A. T. Gallant, P. Hunt, B. Kootte, Y. Lan, S. F. Paul, M. L. Phan, M. P. Reiter, B. Schultz, D. Short, C. Andreoiu, M. Brodeur, I. Dillmann, G. Gwinner, A. A. Kwiatkowski, K. G. Leach, and J. Dilling, Phys. Rev. C 97, 024312 (2018).
- Nesterenko et al. (2020) D. Nesterenko, A. Kankainen, J. Kostensalo, C. Nobs, A. Bruce, O. Beliuskina, L. Canete, T. Eronen, E. Gamba, S. Geldhof, R. de Groote, A. Jokinen, J. Kurpeta, I. Moore, L. Morrison, Z. Podolyák, I. Pohjalainen, S. Rinta-Antila, A. de Roubin, M. Rudigier, J. Suhonen, M. Vilén, V. Virtanen, and J. Äystö, Physics Letters B 808, 135642 (2020).
- Manea et al. (2020) V. Manea, J. Karthein, D. Atanasov, M. Bender, K. Blaum, T. E. Cocolios, S. Eliseev, A. Herlert, J. D. Holt, W. J. Huang, Y. A. Litvinov, D. Lunney, J. Menéndez, M. Mougeot, D. Neidherr, L. Schweikhard, A. Schwenk, J. Simonis, A. Welker, F. Wienholtz, and K. Zuber, Phys. Rev. Lett. 124, 092502 (2020).
- Izzo et al. (2021) C. Izzo, J. Bergmann, K. A. Dietrich, E. Dunling, D. Fusco, A. Jacobs, B. Kootte, G. Kripkó-Koncz, Y. Lan, E. Leistenschneider, E. M. Lykiardopoulou, I. Mukul, S. F. Paul, M. P. Reiter, J. L. Tracy, C. Andreoiu, T. Brunner, T. Dickel, J. Dilling, I. Dillmann, G. Gwinner, D. Lascar, K. G. Leach, W. R. Plaß, C. Scheidenberger, M. E. Wieser, and A. A. Kwiatkowski, Phys. Rev. C 103, 025811 (2021).
- Mougeot et al. (2021) M. Mougeot, D. Atanasov, J. Karthein, R. N. Wolf, P. Ascher, K. Blaum, K. Chrysalidis, G. Hagen, J. D. Holt, W. J. Huang, G. R. Jansen, I. Kulikov, Y. A. Litvinov, D. Lunney, V. Manea, T. Miyagi, T. Papenbrock, L. Schweikhard, A. Schwenk, T. Steinsberger, S. R. Stroberg, Z. H. Sun, A. Welker, F. Wienholtz, S. G. Wilkins, and K. Zuber, Nature Physics 17, 1099 (2021).
- Nesterenko et al. (2023) D. A. Nesterenko, J. Ruotsalainen, M. Stryjczyk, A. Kankainen, L. Al Ayoubi, O. Beliuskina, P. Delahaye, T. Eronen, M. Flayol, Z. Ge, W. Gins, M. Hukkanen, A. Jaries, D. Kahl, D. Kumar, S. Nikas, A. Ortiz-Cortes, H. Penttilä, D. Pitman-Weymouth, A. Raggio, M. Ramalho, M. Reponen, S. Rinta-Antila, J. Romero, A. de Roubin, P. C. Srivastava, J. Suhonen, V. Virtanen, and A. Zadvornaya, Phys. Rev. C 108, 054301 (2023).
- Nies et al. (2023) L. Nies, D. Atanasov, M. Athanasakis-Kaklamanakis, M. Au, K. Blaum, J. Dobaczewski, B. S. Hu, J. D. Holt, J. Karthein, I. Kulikov, Y. A. Litvinov, D. Lunney, V. Manea, T. Miyagi, M. Mougeot, L. Schweikhard, A. Schwenk, K. Sieja, and F. Wienholtz, Phys. Rev. Lett. 131, 022502 (2023).
- Jaries et al. (2023) A. Jaries, M. Stryjczyk, A. Kankainen, L. Al Ayoubi, O. Beliuskina, P. Delahaye, T. Eronen, M. Flayol, Z. Ge, W. Gins, M. Hukkanen, D. Kahl, S. Kujanpää, D. Kumar, I. D. Moore, M. Mougeot, D. A. Nesterenko, S. Nikas, H. Penttilä, D. Pitman-Weymouth, I. Pohjalainen, A. Raggio, W. Rattanasakuldilok, A. de Roubin, J. Ruotsalainen, and V. Virtanen, Phys. Rev. C 108, 064302 (2023).
- Ge et al. (2024) Z. Ge, M. Reponen, T. Eronen, B. Hu, M. Kortelainen, A. Kankainen, I. Moore, D. Nesterenko, C. Yuan, O. Beliuskina, L. Cañete, R. de Groote, C. Delafosse, T. Dickel, A. de Roubin, S. Geldhof, W. Gins, J. D. Holt, M. Hukkanen, A. Jaries, A. Jokinen, A. Koszorús, G. Kripkó-Koncz, S. Kujanpää, Y. H. Lam, S. Nikas, A. Ortiz-Cortes, H. Penttilä, D. Pitman-Weymouth, W. Plaß, I. Pohjalainen, A. Raggio, S. Rinta-Antila, J. Romero, M. Stryjczyk, M. Vilen, V. Virtanen, and A. Zadvornaya, Phys. Rev. Lett. 133, 132503 (2024).
- Jaries et al. (2024) A. Jaries, M. Stryjczyk, A. Kankainen, L. Al Ayoubi, O. Beliuskina, L. Canete, R. P. de Groote, C. Delafosse, P. Delahaye, T. Eronen, M. Flayol, Z. Ge, S. Geldhof, W. Gins, M. Hukkanen, P. Imgram, D. Kahl, J. Kostensalo, S. Kujanpää, D. Kumar, I. D. Moore, M. Mougeot, D. A. Nesterenko, S. Nikas, D. Patel, H. Penttilä, D. Pitman-Weymouth, I. Pohjalainen, A. Raggio, M. Ramalho, M. Reponen, S. Rinta-Antila, A. de Roubin, J. Ruotsalainen, P. C. Srivastava, J. Suhonen, M. Vilen, V. Virtanen, and A. Zadvornaya, Phys. Rev. C 110, 034326 (2024).
- Ruotsalainen et al. (2024) J. Ruotsalainen, D. A. Nesterenko, M. Stryjczyk, A. Kankainen, L. A. Ayoubi, O. Beliuskina, L. Canete, P. Chauveau, R. P. de Groote, P. Delahaye, T. Eronen, M. Flayol, Z. Ge, S. Geldhof, W. Gins, M. Hukkanen, A. Jaries, D. Kahl, D. Kumar, I. D. Moore, S. Nikas, H. Penttilä, D. Pitman-Weymouth, A. Raggio, S. Rinta-Antila, A. de Roubin, M. Vilen, V. Virtanen, and M. Winter, “High-precision mass measurements of the ground and isomeric states in 124,125Ag,” (2024), arXiv:2408.14181 [nucl-ex] .
- Yordanov et al. (2013) D. T. Yordanov, D. L. Balabanski, J. Bieroń, M. L. Bissell, K. Blaum, I. Budinčević, S. Fritzsche, N. Frömmgen, G. Georgiev, C. Geppert, M. Hammen, M. Kowalska, K. Kreim, A. Krieger, R. Neugart, W. Nörtershäuser, J. Papuga, and S. Schmidt, Phys. Rev. Lett. 110, 192501 (2013).
- Ferrer et al. (2014) R. Ferrer, N. Bree, T. Cocolios, I. Darby, H. De Witte, W. Dexters, J. Diriken, J. Elseviers, S. Franchoo, M. Huyse, N. Kesteloot, Y. Kudryavtsev, D. Pauwels, D. Radulov, T. Roger, H. Savajols, P. Van Duppen, and M. Venhart, Physics Letters B 728, 191 (2014).
- Hammen et al. (2018) M. Hammen, W. Nörtershäuser, D. L. Balabanski, M. L. Bissell, K. Blaum, I. Budinčević, B. Cheal, K. T. Flanagan, N. Frömmgen, G. Georgiev, C. Geppert, M. Kowalska, K. Kreim, A. Krieger, W. Nazarewicz, R. Neugart, G. Neyens, J. Papuga, P.-G. Reinhard, M. M. Rajabali, S. Schmidt, and D. T. Yordanov, Phys. Rev. Lett. 121, 102501 (2018).
- Gorges et al. (2019) C. Gorges, L. V. Rodríguez, D. L. Balabanski, M. L. Bissell, K. Blaum, B. Cheal, R. F. Garcia Ruiz, G. Georgiev, W. Gins, H. Heylen, A. Kanellakopoulos, S. Kaufmann, M. Kowalska, V. Lagaki, S. Lechner, B. Maaß, S. Malbrunot-Ettenauer, W. Nazarewicz, R. Neugart, G. Neyens, W. Nörtershäuser, P.-G. Reinhard, S. Sailer, R. Sánchez, S. Schmidt, L. Wehner, C. Wraith, L. Xie, Z. Y. Xu, X. F. Yang, and D. T. Yordanov, Phys. Rev. Lett. 122, 192502 (2019).
- Yordanov et al. (2020) D. T. Yordanov, L. V. Rodríguez, D. L. Balabanski, J. Bieroń, M. L. Bissell, K. Blaum, B. Cheal, J. Ekman, G. Gaigalas, R. F. Garcia Ruiz, G. Georgiev, W. Gins, M. R. Godefroid, C. Gorges, Z. Harman, H. Heylen, P. Jönsson, A. Kanellakopoulos, S. Kaufmann, C. H. Keitel, V. Lagaki, S. Lechner, B. Maaß, S. Malbrunot-Ettenauer, W. Nazarewicz, R. Neugart, G. Neyens, W. Nörtershäuser, N. S. Oreshkina, A. Papoulia, P. Pyykkö, P.-G. Reinhard, S. Sailer, R. Sánchez, S. Schiffmann, S. Schmidt, L. Wehner, C. Wraith, L. Xie, Z. Xu, and X. Yang, Communications Physics 3, 107 (2020).
- Reponen et al. (2021) M. Reponen, R. P. de Groote, L. Al Ayoubi, O. Beliuskina, M. L. Bissell, P. Campbell, L. Cañete, B. Cheal, K. Chrysalidis, C. Delafosse, A. de Roubin, C. S. Devlin, T. Eronen, R. F. Garcia Ruiz, S. Geldhof, W. Gins, M. Hukkanen, P. Imgram, A. Kankainen, M. Kortelainen, Á. Koszorús, S. Kujanpää, R. Mathieson, D. A. Nesterenko, I. Pohjalainen, M. Vilén, A. Zadvornaya, and I. D. Moore, Nature Communications 12, 4596 (2021).
- Geldhof et al. (2022) S. Geldhof, M. Kortelainen, O. Beliuskina, P. Campbell, L. Caceres, L. Cañete, B. Cheal, K. Chrysalidis, C. S. Devlin, R. P. de Groote, A. de Roubin, T. Eronen, Z. Ge, W. Gins, A. Koszorus, S. Kujanpää, D. Nesterenko, A. Ortiz-Cortes, I. Pohjalainen, I. D. Moore, A. Raggio, M. Reponen, J. Romero, and F. Sommer, Phys. Rev. Lett. 128, 152501 (2022).
- Vernon et al. (2022) A. R. Vernon, R. F. Garcia Ruiz, T. Miyagi, C. L. Binnersley, J. Billowes, M. L. Bissell, J. Bonnard, T. E. Cocolios, J. Dobaczewski, G. J. Farooq-Smith, K. T. Flanagan, G. Georgiev, W. Gins, R. P. de Groote, R. Heinke, J. D. Holt, J. Hustings, Á. Koszorús, D. Leimbach, K. M. Lynch, G. Neyens, S. R. Stroberg, S. G. Wilkins, X. F. Yang, and D. T. Yordanov, Nature 607, 260 (2022).
- Karthein et al. (2024) J. Karthein, C. M. Ricketts, R. F. Garcia Ruiz, J. Billowes, C. L. Binnersley, T. E. Cocolios, J. Dobaczewski, G. J. Farooq-Smith, K. T. Flanagan, G. Georgiev, W. Gins, R. P. de Groote, F. P. Gustafsson, J. D. Holt, A. Kanellakopoulos, Á. Koszorús, D. Leimbach, K. M. Lynch, T. Miyagi, W. Nazarewicz, G. Neyens, P.-G. Reinhard, B. K. Sahoo, A. R. Vernon, S. G. Wilkins, X. F. Yang, and D. T. Yordanov, Nature Physics (2024), 10.1038/s41567-024-02612-y.
- de Groote et al. (2024) R. de Groote, D. Nesterenko, A. Kankainen, M. Bissell, O. Beliuskina, J. Bonnard, P. Campbell, L. Canete, B. Cheal, C. Delafosse, A. de Roubin, C. Devlin, J. Dobaczewski, T. Eronen, R. Garcia Ruiz, S. Geldhof, W. Gins, M. Hukkanen, P. Imgram, R. Mathieson, A. Koszorús, I. Moore, I. Pohjalainen, M. Reponen, B. van den Borne, M. Vilén, and S. Zadvornaya, Physics Letters B 848, 138352 (2024).
- Eliseev et al. (2013) S. Eliseev, K. Blaum, M. Block, C. Droese, M. Goncharov, E. Minaya Ramirez, D. A. Nesterenko, Y. N. Novikov, and L. Schweikhard, Phys. Rev. Lett. 110, 082501 (2013).
- Eliseev et al. (2014) S. Eliseev, K. Blaum, M. Block, A. Dörr, C. Droese, T. Eronen, M. Goncharov, M. Höcker, J. Ketter, E. M. Ramirez, D. A. Nesterenko, Y. N. Novikov, and L. Schweikhard, Applied Physics B 114, 107 (2014).
- Nesterenko et al. (2018) D. A. Nesterenko, T. Eronen, A. Kankainen, L. Canete, A. Jokinen, I. D. Moore, H. Penttilä, S. Rinta-Antila, A. de Roubin, and M. Vilen, The European Physical Journal A 54, 154 (2018).
- Koponen et al. (1989) V. Koponen, J. Äystö, J. Honkanen, P. Jauho, H. Penttilä, J. Suhonen, P. Taskinen, K. Rykaczewski, J. Żylicz, and C. N. Davids, Zeitschrift fur Physik A Hadrons and Nuclei 333, 339 (1989).
- Batchelder et al. (2005) J. C. Batchelder, J.-C. Bilheux, C. R. Bingham, H. K. Carter, D. Fong, P. E. Garrett, R. Grzywacz, J. H. Hamilton, D. J. Hartley, J. K. Hwang, W. Krolas, W. D. Kulp, Y. Larochelle, A. Piechaczek, A. V. Ramayya, K. P. Rykaczewski, E. H. Spejewski, D. W. Stracener, M. N. Tantawy, J. A. Winger, J. L. Wood, and E. F. Zganjar, Phys. Rev. C 72, 044306 (2005).
- Batchelder et al. (2012) J. C. Batchelder, N. T. Brewer, R. E. Goans, R. Grzywacz, B. O. Griffith, C. Jost, A. Korgul, S. H. Liu, S. V. Paulauskas, E. H. Spejewski, and D. W. Stracener, Phys. Rev. C 86, 064311 (2012).
- Janas et al. (1993) Z. Janas, J. Äystö, K. Eskola, P. Jauho, A. Jokinen, J. Kownacki, M. Leino, J. Parmonen, H. Penttilä, J. Szerypo, and J. Żylicz, Nuclear Physics A 552, 340 (1993).
- Wang et al. (2003) Y. Wang, S. Rinta-Antila, P. Dendooven, J. Huikari, A. Jokinen, V. S. Kolhinen, G. Lhersonneau, A. Nieminen, S. Nummela, H. Penttilä, K. Peräjärvi, J. Szerypo, J. C. Wang, and J. Äystö, Phys. Rev. C 67, 064303 (2003).
- Rissanen et al. (2007) J. Rissanen, V. V. Elomaa, T. Eronen, J. Hakala, A. Jokinen, S. Rahaman, S. Rinta-Antila, and J. Äystö, The European Physical Journal A 34, 113 (2007).
- Karvonen et al. (2008) P. Karvonen, I. Moore, T. Sonoda, T. Kessler, H. Penttilä, K. Peräjärvi, P. Ronkanen, and J. Äystö, Nuclear Instruments and Methods in Physics Research Section B: Beam Interactions with Materials and Atoms 266, 4794 (2008).
- Nieminen et al. (2001) A. Nieminen, J. Huikari, A. Jokinen, J. Äystö, P. Campbell, and E. Cochrane, Nucl. Instrum. Meth. Phys. Res. A 469, 244 (2001).
- Eronen et al. (2012) T. Eronen, V. S. Kolhinen, V. V. Elomaa, D. Gorelov, U. Hager, J. Hakala, A. Jokinen, A. Kankainen, P. Karvonen, S. Kopecky, I. D. Moore, H. Penttilä, S. Rahaman, S. Rinta-Antila, J. Rissanen, A. Saastamoinen, J. Szerypo, C. Weber, and J. Äystö, The European Physical Journal A 48, 46 (2012).
- de Groote et al. (2020) R. de Groote, A. de Roubin, P. Campbell, B. Cheal, C. Devlin, T. Eronen, S. Geldhof, I. Moore, M. Reponen, S. Rinta-Antila, and M. Schuh, Nuclear Instruments and Methods in Physics Research Section B: Beam Interactions with Materials and Atoms 463, 437 (2020).
- Savard et al. (1991) G. Savard, S. Becker, G. Bollen, H. J. Kluge, R. B. Moore, T. Otto, L. Schweikhard, H. Stolzenberg, and U. Wiess, Phys. Lett. A 158, 247 (1991).
- Eronen et al. (2008) T. Eronen, V.-V. Elomaa, U. Hager, J. Hakala, A. Jokinen, A. Kankainen, S. Rahaman, J. Rissanen, C. Weber, and J. Äystö, Nucl. Instrum. Meth. Phys. Res. B 266, 4527 (2008).
- Wang et al. (2021) M. Wang, W. Huang, F. Kondev, G. Audi, and S. Naimi, Chinese Physics C 45, 030003 (2021).
- Vilén et al. (2020) M. Vilén, L. Canete, B. Cheal, A. Giatzoglou, R. de Groote, A. de Roubin, T. Eronen, S. Geldhof, A. Jokinen, A. Kankainen, I. Moore, D. Nesterenko, H. Penttilä, I. Pohjalainen, M. Reponen, and S. Rinta-Antila, Nuclear Instruments and Methods in Physics Research Section B: Beam Interactions with Materials and Atoms 463, 382 (2020).
- Nesterenko et al. (2021) D. Nesterenko, T. Eronen, Z. Ge, A. Kankainen, and M. Vilen, The European Physical Journal A 57, 302 (2021).
- Roux et al. (2013) C. Roux, K. Blaum, M. Block, C. Droese, S. Eliseev, M. Goncharov, F. Herfurth, E. M. Ramirez, D. A. Nesterenko, Y. N. Novikov, and L. Schweikhard, The European Physical Journal D 67, 146 (2013).
- Gins et al. (2018) W. Gins, R. P. de Groote, M. L. Bissell, C. G. Buitrago, R. Ferrer, K. M. Lynch, G. Neyens, and S. Sels, Computer Physics Communications 222, 286 (2018).
- Cannarozzo et al. (2023) S. Cannarozzo, S. Pomp, A. Solders, A. Al-Adili, A. Göök, and A. Koning, The European Physical Journal A 59, 295 (2023).
- Persson (2013) J. Persson, Atomic Data and Nuclear Data Tables 99, 62 (2013).
- Dahmen and Penselin (1967) H. Dahmen and S. Penselin, Zeitschrift für Physik 200, 456 (1967).
- Sahm and Schwenk (1974) W. Sahm and A. Schwenk, Zeitschrift für Naturforschung A 29, 1763 (1974).
- Mertzimekis et al. (2016) T. Mertzimekis, K. Stamou, and A. Psaltis, Nuclear Instruments and Methods in Physics Research Section A: Accelerators, Spectrometers, Detectors and Associated Equipment 807, 56 (2016).
- Fischer et al. (1975) W. Fischer, H. Hühnermann, and T. Meier, Zeitschrift für Physik A Atoms and Nuclei 274, 79 (1975).
- Berkes et al. (1984) I. Berkes, B. Hlimi, G. Marest, E. Sayouty, R. Coussement, F. Hardeman, P. Put, and G. Scheveneels, Physical Review C 30, 2026 (1984).
- Ohayon et al. (2024) B. Ohayon, J. E. Padilla-Castillo, S. C. Wright, G. Meijer, and B. K. Sahoo, Phys. Rev. Res. 6, 033040 (2024).
- Penttilä et al. (1990) H. Penttilä, J. Äystö, P. Jauho, A. Jokinen, J. M. Parmonen, P. Taskinen, K. Eskola, M. Leino, P. Dendooven, and C. N. Davids, Physica Scripta T32, 38 (1990).
- Kondev et al. (2021) F. Kondev, M. Wang, W. Huang, S. Naimi, and G. Audi, Chinese Physics C 45, 030001 (2021).
- Batchelder et al. (2009) J. C. Batchelder, J. L. Wood, P. E. Garrett, K. L. Green, K. P. Rykaczewski, J. C. Bilheux, C. R. Bingham, H. K. Carter, D. Fong, R. Grzywacz, J. H. Hamilton, D. J. Hartley, J. K. Hwang, W. Krolas, W. D. Kulp, Y. Larochelle, A. Piechaczek, A. V. Ramayya, E. H. Spejewski, D. W. Stracener, M. N. Tantawy, J. A. Winger, and E. F. Zganjar, Phys. Rev. C 80, 054318 (2009).
- Porquet et al. (2003) M.-G. Porquet, T. Venkova, A. Astier, A. Bauchet, I. Deloncle, N. Buforn, L. Donadille, O. Dorvaux, B. J. P. Gall, S. Lalkovski, R. L. as, M. Meyer, A. Minkova, A. Prévost, N. Redon, N. Schulz, and O. Stézowski, The European Physical Journal A - Hadrons and Nuclei 18, 25 (2003).
- Kim et al. (2017) Y. Kim, S. Biswas, M. Rejmund, A. Navin, A. Lemasson, S. Bhattacharyya, M. Caamaño, E. Clément, G. de France, and B. Jacquot, Physics Letters B 772, 403 (2017).
- Fogelberg et al. (1971) B. Fogelberg, A. Bäcklin, and T. Nagarajan, Physics Letters B 36, 334 (1971).
- Wang et al. (2017) E. H. Wang, J. H. Hamilton, A. V. Ramayya, Y. X. Liu, H. J. Li, A. C. Dai, W. Y. Liang, F. R. Xu, J. K. Hwang, S. H. Liu, N. T. Brewer, Y. X. Luo, J. O. Rasmussen, Y. Sun, S. J. Zhu, G. M. Ter-Akopian, and Y. T. Oganessian, Phys. Rev. C 95, 064311 (2017).
- Hill (1978) J. C. Hill, BNL-50847; CONF-771078- , 169 (1978).
- Fritzs and Griffiths (1967) K. Fritzs and K. Griffiths, Radiochimica Acta 7, 59 (1967).
- Weiss et al. (1968) H. V. Weiss, J. M. Fresco, and W. L. Reichert, Phys. Rev. 172, 1266 (1968).
- Heyde (1994) K. L. G. Heyde, “The nuclear shell model,” in The Nuclear Shell Model: Study Edition (Springer Berlin Heidelberg, Berlin, Heidelberg, 1994) pp. 58–154.
- Winnacker et al. (1976) A. Winnacker, H. Ackermann, D. Dubbers, M. Grupp, P. Heitjans, and H.-J. Stöck-Mann, Nuclear Physics A 261, 261 (1976).
- Golovko et al. (2010) V. V. Golovko, I. S. Kraev, T. Phalet, B. Delauré, M. Beck, V. Y. Kozlov, S. Coeck, F. Wauters, P. Herzog, C. Tramm, D. Zákoucký, D. Vénos, D. Srnka, M. Honusek, U. Köster, and N. Severijns, Phys. Rev. C 81, 054323 (2010).
- Greenebaum and Phillips (1974) B. Greenebaum and E. A. Phillips, Phys. Rev. C 9, 2028 (1974).
- Eder et al. (1985) R. Eder, E. Hagn, and E. Zech, Phys. Rev. C 31, 190 (1985).
- Dinger et al. (1989) U. Dinger, J. Eberz, G. Huber, R. Menges, R. Kirchner, O. Klepper, T. Kühl, and D. Marx, Nuclear Physics A 503, 331 (1989).
- Spence and McDermott (1972) P. Spence and M. McDermott, Physics Letters A 42, 273 (1972).
- Eberz et al. (1987a) J. Eberz, U. Dinger, G. Huber, H. Lochmann, R. Menges, R. Neugart, R. Kirchner, O. Klepper, T. Kühl, D. Marx, G. Ulm, and K. Wendt, Nuclear Physics A 464, 9 (1987a).
- Eberz et al. (1987b) J. Eberz, U. Dinger, G. Huber, H. Lochmann, R. Menges, G. Ulm, R. Kirchner, O. Klepper, T. Kühl, and D. Marx, Zeitschrift für Physik A Atomic Nuclei 326, 121 (1987b).
- Chaney and McDermott (1969) R. Chaney and M. McDermott, Physics Letters A 29, 103 (1969).
- Proctor (1950) W. G. Proctor, Phys. Rev. 79, 35 (1950).
- Ortiz Cortes (2023) A. Ortiz Cortes, Palladium: a Study of Nuclear Deformation of a Refractory Element. Laser Spectroscopy at the IGISOL, Ph.D. thesis, University of Jyväskylä (2023), http://urn.fi/URN:ISBN:978-951-39-9287-3.
- Anselment et al. (1986) M. Anselment, K. Bekk, A. Hanser, H. Hoeffgen, G. Meisel, S. Göring, H. Rebel, and G. Schatz, Physical Review C 34, 1052 (1986).
- Piller et al. (1990) C. Piller, C. Gugler, R. Jacot-Guillarmod, L. Schaller, L. Schellenberg, H. Schneuwly, G. Fricke, T. Hennemann, and J. Herberz, Physical Review C 42, 182 (1990).
- Dimmling et al. (1975) F. Dimmling, D. Riegel, K.-G. Rensfelt, and C.-J. Herrlander, Physics Letters B 55, 293 (1975).
- Le Blanc et al. (2004) F. Le Blanc, L. Cabaret, E. Cottereau, J. Crawford, S. Essabaa, J. Genevey, R. Horn, G. Huber, J. Lassen, J. Lee, G. Le Scornet, J. Obert, J. Oms, A. Ouchrif, J. Pinard, B. Roussie‘re, J. Sauvage, and D. Verney, Nuclear Physics A 734, 437 (2004).
- de Groote et al. (2020) R. P. de Groote, M. L. Bissel, T. E. Cocolios, B. S. Cooper, K. Flanagan, S. Franchoo, R. F. G. Ruiz, R. Heinke, J. Johnson, A. Koszorús, Y. C. Liu, G. Neyens, F. P. Gustafsson, C. M. Ricketts, M. Reponen, S. W. Bai, A. R. Vernon, S. G. Wilkins, and X. F. Yang, Collinear resonance ionization spectroscopy of silver between N=50 and N=82, Tech. Rep. (CERN, Geneva, 2020).
- Cocolios (2020) T. E. Cocolios, Absolute charge radii of radioactive isotopes measured by muonic x ray spectroscopy at PSI, Tech. Rep. (CERN, Geneva, 2020).
Supplementary Material: Binding energies, charge radii, spins and moments: odd-odd Ag isotopes and discovery of a new isomer
V Additional results on odd-even Ag
Results obtained from the same experimental campaign as this work on odd-even 113-121Ag have been partially published before in Refs. Reponen et al. (2021); de Groote et al. (2024). Here we provide the -parameters, isotope shifts, and changes in mean-squared radii not published before (see Table 3).
| Nuclide | A(P3/2) | r | ||
| (MHz) | (MHz) | (fm2) | ||
| 113Agm | 173.7(13) | 0.3147(14)[45] | ||
| 115Agm | 173.5(7) | 0.446(2)[6] | ||
| 117Agm | 172.5(4) | 0.561(5)[8] | ||
| 119Agm | 177.2(5) | 0.669(7)[10] | ||
| 121Ag | 172.5(16) | 0.7570(14)[110] |
VI Extraction of relative production yields from hyperfine spectra
The relative production yields were approximated by comparing the integration of hyperfine peaks, i.e. transitions, for different nuclear states. First, the background was subtracted from every spectrum. Then, which is the sum of the integrals of all hyperfine transitions for all hyperfine spectra of a nuclear state , is calculated with Eq. 8. is corrected for the acquisition time and for Racah intensity via the NRacahh factor defined in Eq. 9. The idea behind Eq. 8 is that nuclear states with the same production rate will result in the same insensitive to the nuclear spin I, the chosen transition h, and the acquisition time i. The relative yield was calculated using Eq. 10.
The integralhi is the integral of the transition h in the spectrum i, is the total number of scans of a nuclear state , NRacahh is the normalized Racah intensity of the transition h and timei is the acquisition time of the spectrum i. Then to retrieve the relative yield of a nuclear state I Eq. 10 was used, where the sum is over the spins of all measured nuclear states. In total, there were four hyperfine spectra measured and eight different sets of transitions were evaluated. The statistical error on each set is derived by calculating the standard deviation between for the different spectra i.
These results are shown in Table 4 together with the ratio for the assigned level order. We note there is some spread on the obtained results in Table 4, likely due to the Racah intensity approximation used in the calculation.
| Chosen hyperfine transitions to integrate | Relative production yield (%) | Ratio | |||
| Spin | |||||
| 8.9(10) | 30.8(35) | 60(4) | 0.45(6) | ||
| 10.6(10) | 36.5(35) | 53(4) | 0.57(7) | ||
| 9.3(10) | 32.1(35) | 59(4) | 0.47(6) | ||
| 11.1(10) | 38.3(33) | 50.6(33) | 0.62(6) | ||
| 9.5(11) | 26.4(29) | 64.0(35) | 0.36(4) | ||
| 9.3(10) | 28.1(34) | 63(4) | 0.39(5) | ||
| 9.9(11) | 23.5(26) | 66.6(34) | 0.31(4) | ||
| 12.6(11) | 30.0(24) | 57.4(29) | 0.42(4) | ||
| Average | 10.2(4) | 30.7(11) | 59.1(13) | 0.443(18) | |